in the initial concentration if [N2O5] is 4.0 M, what is the concentration after 155 seconds
1
Expert's answer
2018-02-24T07:17:07-0500
The decomposition of dinitrogen pentoxide (N2O5) has a rate constant of 4.50 x 10-5 s-1 at 300 K*.

k = 1/tau ln C_0/C

4.50 x 10-5 s-1 = 1 / 155 s ln 4.0 M / C
ln 4.0 M / C = 0.006975
4 M / C = 1.01
C = 3.96 M

Answer: 3.96 M.